OpenStudy (anonymous):

the curve y = 3x^{1/4}. The shaded region is bounded by the curve, the x-axis and the lines x = 1 and x = 4. Find the volume of the solid obtained when this shaded region is rotated completely about the x-axis, giving your answer in terms of π.

OpenStudy (dumbcow):

\[V = \pi \int\limits_{1}^{4}(3x^{1/4})^{2} dx\]

OpenStudy (anonymous):

what is (x^1/4)^2?

OpenStudy (dumbcow):

just multiply the exponents --> (x^1/4)^2 = x^2/4 = x^1/2

OpenStudy (anonymous):

Why wouldn't it be 1/16?

OpenStudy (anonymous):

u there? need help?

OpenStudy (anonymous):

I understand everything apart from why 1/4^2 is 1/2

OpenStudy (anonymous):